Name MULTIPLE CHOICE. Choose the one alternative that best completes the statement or answens Chem 1A Weekly Quiz5 1) Determine the total the question. pressure of all gases (at STP) formed when 500mLofNT(C3H5NOI.d. 1.60 g/ml, molar mass 227.10 g/mol) reacts according to the following reaction. 1)_ - 4 C3H5(NO3130)-6 N2(8)+02) 12 CO2(g) 10 H2og) A) 57.2L B) 29.6L ) 175L D) 493L E) 448 L at volume (L) of 0.250 M HNO3 is nequired to neutralize a solution peepared by dissolving 175 2) g of NaOH in 350 mL of water? A) 0.44 B) 1.75 C) 0070 D) 50.0 E) 1.75- 10-3 3) 3) An ideal gas differs from a real gas in that the molecules of an ideal gas A) have appreciable molecular volumes B) have no kinetic energy C) have an average molecular mass D) have a molecular weight of zero E) have no attraction for one another 4) Arrange the following gases in order of increasing average molecular speed at 25 C He, 02 CO2 N2 5 5) A sample of an unknown volatile liquid was injected into a Dumas flask (mglask- 27 0928 g lask 0.1040 L) and heated until no visible traces of the liquid could be found. The flask and its contents were then rapidly cooled and reweighed (mlaskevapor 27 4593 g). The atmospheric pressure and temperature during the experiment were 0.976 atm and 180 ? respectively. The unknown volatile liquidwas_ B) C6H14 C) C6H6 D) C6H12 E) C7H14 A) C7H16 6 6) A large balloon is initially filled to a volume of 25.0 L at 353 K and a pressure of 2575 mm Hg What volume of gas will the balloon contain at 1.35 atm and 253 K? E) 22.2L B) 11.4L C) 45.0 L D) 58.6L A) 87.5L
Answer #1

1.

Mass of 50.0 mL of TNT = 50.0 mL * 1.60 g/mL = 80.0 g

Moles of TNT = (80.0 g) / (227.10 g/mol) = 0.35 moles

Now,
4 C3H5(NO3)3(l) --> 6 N2(g) + O2(g) + 12 CO2(g) + 10 H2O(g)

From the above reaction equation, 4 moles of TNT produces, a total of 6+1+12+10 = 29 moles of gases.

So, 0.35 moles of TNT will produce 0.35*(29/4) = 2.54 moles of gases.

From the law of gases

PV = nRT
where P the pressure, V the Volume, n the number of moles, R the gas constant and T the absolute temperature.

At STP, P = 1 atm, T = 273 K

At the moment of explosion the gases take up the same volume as that of the TNT (50.0 mL = 0.050 L).

So,

P * (0.050 L) = (2.54 mol) * (0.082 L atm mol-1 K-1) * (273 K)
or, P = [(2.54 mol) * (0.082 L atm mol-1 K-1) * (273 K) ] / (0.050 L)
         = 1137.2 atm

2.

Mass of NaOH = 17.5 g
Molar mass of NaOH = 40 g/mol
So, moles of NaOH = (17.5 g) / (40 g/mol)
                                  = 0.4375 moles

Since, HNO3 and NaOH reacts in 1:1 ratio. So, moles of HNO3 = 0.4375 moles

Molarity of HNO3 = 0.250 M

So, volume of HNO3 = (0.4375 mol) / (0.250 M) = 1.75 L

Answer is option (B) 1.75 L

?3.

Answer is option (E) have no attraction for one another

4.

The molecular speed of gases at constant temperature varies inversely with the square root of the molecular mass. So the lightest molecule will go the fastest.
The actual order would be CO2 < O2 < N2 < He.

Answer is option (A).

6.

We know that

P1V1/T1 = P2V2/T2

or, (2575 mm Hg) x (25.0 L) / (353 K) = (1.35 atm) V2 / (253 K)

or, (2575 x 0.00132 atm) x (25.0 L) / (353 K) = (1.35 atm) V2 / (253 K)

or, (3.40 atm) x (25.0 L) / (353 K) = (1.35 atm) V2 / (253 K)

or, V2 = (3.40 atm) x (25.0 L) x (253 K) / [ (353 K) x (1.35 atm) ]

or, V2 = 45.1 L

Answer is option (B) 45.0 L